### v1

# EgoBlind: Towards Egocentric Visual Assistance for the Blind People

Abstract
We present EgoBlind, the first egocentric VideoQA dataset collected from blind individuals to evaluate the assistive capabilities of contemporary multimodal large language models (MLLMs). EgoBlind comprises 1,210 videos that record the daily lives of real blind users from a first-person perspective. It also features 4,927 questions directly posed or generated and verified by blind individuals to reflect their needs for visual assistance under various scenarios. We provide each question with an average of 3 reference answers to alleviate subjective evaluation. Using EgoBlind, we comprehensively evaluate 15 leading MLLMs and find that all models struggle, with the best performers achieving accuracy around 56%, far behind human performance of 87.4%. To guide future advancements, we identify and summarize major limitations of existing MLLMs in egocentric visual assistance for the blind and provide heuristic suggestions for improvement. With these efforts, we hope EgoBlind can serve as a valuable foundation for developing more effective AI assistants to enhance the independence of the blind individuals’ lives.
1 Introduction
The rapid advancement of multimodal large language models (MLLMs) [1, 20, 26, 35, 29, 34] has significantly improved the performance of visual question answering (VQA). However, existing VQA datasets primarily focus on third-person perspective [12, 37] or general-purpose image and video understanding [17, 6, 28], overlooking critical real-world applications such as visual assistance for the visually-impaired individuals who constitute an overwhelming population of 2.2 billion people around the world [30, 14]. The research is especially scarce in assisting the blind from a first-person perspective [33].
Figure 1: Real example of EgoBlind about blind user seeking egocentric visual assistance. The first time when the user puts their hand on the knob of the microwave, they are curious about what the knob control. The second time after they tune the knob, they are interested to know the position of the knob. The third time when they put their hand on another knob, they are curious about what knob it is. Finally, they want to know if it is correct to turn the knob clock-wise, to set the cooking time.
In light of this, we construct EgoBlind – the first egocentric VideoQA dataset designed to benchmark and advance MLLMs towards live egocentric visual assistance for blind people. EgoBlind comprises 1,210 egocentric videos that capture the visual experiences of blind individuals and 4,927 questions that are directly posed or automatically generated and further verified by blind users to reflect their assistive needs in exploring their surroundings. To categorize these needs for better analysis, the questions are grouped into six key types: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. We set the QA task as real-time (timestamp restricted) and open-form answer generation to better align with its live assistance nature. We provide multiple reference answers for each question for effective evaluation and carefully prompt the LLM for judgment.
By focusing on blind users’ needs for real-time egocentric visual assistance, we aim to explore several key challenges in video-language learning, including egocentric dynamic scene understanding with poor visual quality (e.g., unstable motion, object blur, and occlusions), real-time context-aware user intention reasoning, and assistance-oriented answer generation. Take Fig. 1 as an example; engaging in the user’s activity from a first-person perspective is crucial to understanding and answering the questions. This specification requires capturing the gaze area and hand motion and reasoning the related visual content spatio-temporally. Finally, this should be achieved with poor quality visual inputs, as the regions of interest are often not in focus (e.g., the control panel in Fig. 1).
With EgoBlind, we benchmark 15 recent MLLMs, covering both the open-source models (e.g., InternVL2 [5], Qwen2-VL [35] and LLaVA-OV [19] and VideoLLaMA3 [44]) and the closed-source ones (e.g., GPT-4o [29] and Gemini 2.0 [34]). For open-source models, we specially take into account the models that 1) achieve the state-of-the-art (SOTA) results on common video QA benchmarks [43, 37, 27, 11], and benchmarks emphasizing egocentric understanding [28, 6, 42]. Additionally, we include three SOTA modes (Video-LLaVA [22] and LLaMA-VID [21], VILA [23]) on the image blind QA dataset VizWiz [14].
[…]
  • Models are poor in assisting with egocentric navigation, safety warning, and tool use, indicating limited knowledge in these fields.
  • Models generate objectively accurate answers about visual content, but the answers often fail to meet blind users’ subjective needs for assistance. This result suggests a weakness in reasoning about user intentions within dynamic egocentric visual contexts.
  • Models, especially the open-source ones, are sycophantic [47]; they hallucinate wrong and potentially malicious answers when the blind users ask about non-existent objects in their surroundings due to blindness.
  • Models fail to reason the temporal contexts and localize objects relative to the blind users’ real-time locations.
Our contributions are: 1) we tackle significant challenges to construct the first VideoQA dataset (EgoBlind) to benchmark and promote research toward egocentric visual assistance for the blind, and 2) we comprehensively analyze the behaviors of the leading MLLMs, revealing lots of limitations and possibilities for improvements.
2 Related Work
2.1 VQA for the Blind
Visual question answering (VQA) is the task of answering user questions about images and videos. As one of its most impactful applications, VQA has been proposed to help visually impaired individuals [2]. Yet, so far the vast majority of advancements have been made in general-purpose VQA [12, 40, 37, 43] where the questions are to test the models’ capability of general-purpose visual understanding, with also the images and videos captured from third-person perspective. The closest in aim is VizWiz [14], which collects data from real human-powered VQA systems for helping blind users (e.g., BeMyEye [^1] http://www.bemyeyes.org/). Despite the significance, VizWiz limits in static images and object-centric questions; it fails to cater to the real-time and broader assistive needs of blind users in exploring the dynamic surroundings. Recent ongoing work VIEW-QA [33] proposes to capture the daily challenges faced by visually impaired individuals via using 360-degree egocentric wearable cameras. However, its videos and questions are collected by training actors to simulate the experiences of blind individuals, limiting its ability to reflect the real blind people’s daily lives and their true needs for visual assistance, especially for those who have been blind for a long time.
2.2 Egocentric VQA
Egocentric VQA has earned increasing interest for its significant application value towards embodied assistance [31, 39]. Early effort EgoVQA [10] is small scale and limits its questions in challenging action recognition from first-person perspective with the absence of the camera wearer in the footage. Subsequent advancement EgoTaskQA [17] uses machine-generated questions to evaluate models’ task understanding capabilities. AssistQ [36] is close to our aim for embodied assistance, but it limits to instructional videos for demonstration of tool use. Ever since the release of Ego4D [13], lots of real-world egocentric QA tasks are explored [24, 32, 3, 28, 9, 7, 6, 42]. However, most of they aim at general-purpose egocentric visual understanding, or focus on a specific aspect of assistance, e.g., QAEgo4D [3] for episodic memory and EgoTextVQA [48] for scene text understanding. To our best knowledge, there is so far no existing egocentric VQA dataset specially collected from real blind individuals.
2.3 MLLMs for VQA
MLLMs have significantly improved the landscape of VQA. By integrating visual information into powerful LLMs through lightweight connection modules (e.g., MLP [25]), MLLMs extend the capabilities of LLMs to converse with images and videos naturally like humans. This brings breakthrough over traditional VQA which answers questions by either multi-choice selection [16, 37] or close-vocabulary classification [40, 43]. While most of the MLLMs are developed for VQA of third-person perspective [39, 27, 45, 22, 38, 21, 8, 35, 44], recent advancements [24, 32, 46, 4, 5, 42] specially include ego videos for understanding. Nonetheless, all of they target a general-purpose egocentric visual understanding. In this paper, we will comprehensively examine their capacities towards egocentric visual assistance for the blind users.
3 EgoBlind Dataset
3.1 Dataset Creation
Initially, we have tried to invite blind volunteers to record their daily lives from a first-person perspective and pose questions about their needs for daily visual assistance, similar to the data collection process of Ego4D [13]. However, we find that this approach presents significant challenges regarding data diversity (we can only get contact to a very limited number of blind people), quality (people often sparsely and coarsely ask what surrounds them to complete the data collection task), as well as ethical and privacy concerns. As an alternative, we collect videos from social media platforms and adopt the annotation pipeline as illustrated in Fig. 2. We elaborate on the details below.
Figure 2: Data annotation pipeline.
3.1.1 Video Collection
We crawl 425 long-form egocentric videos from video media platforms such as Bilibili and TikTok, all originating from visually impaired content creators. These videos, captured from the first-person perspective of blind individuals using GoPro or mobile phone, document their daily lives (such as in traveling, cooking, navigating, social gatherings, medical visiting.) and provide real-world data on how visually impaired individuals perform tasks and solve problems in dynamic environments. Since a single long video often contains a vast array of visual content, e.g., a summarized record of a visually impaired creator’s week, we manually parition the videos into distinct segments following the traces of splicing. During editing, we delete video moments with sharp scene transitions and big subtitles to enhance the focus of each clip. Ultimately, we obtain 1,210 video clips with an average duration of 35.8 seconds.
3.1.2 Question Collection
We obtain the questions via three different approaches: 1) Manual Extraction: We manually extract the visual-assistance related questions directly posed by the blind users in the videos. 2) AI-Generation: We prompt GPT-4o [29] to act as blind user and generate questions by engaging in their egocentric perspective. The generated questions are further verified and edited as necessary by both sighted and visually impaired people, ensuring their correctness and alignment with blind users’ needs. 3) Blind User Annotation: For some videos, we introduce their major contents to the blind individuals and invite them to annotate questions when confronted with the visual situations. The actual annotation is conducted within the blind user verification stage. The collection details are as follows:
Table 1: QA examples categorized by different assistance needs.
Category #Q / #V QA Examples (Multiple reference answers)
Information Reading 2283 / 962 Q: What floor is the elevator currently on?
A: 1st floor / At this time on the first floor.
Safety
Warnings
1181 / 629 Q: Is it safe to cross the street now?
A: No. / No, there are cars on both sides. / A car is about to pass through the street, it is recommended to wait a little.
Navigation 669 / 389 Q: Where is the entrance to the building?
A: Just a few steps ahead of you. / Directly in front.
Social
Communication
140 / 119 Q: Who is the person talking to me?
A: He is a delivery guy. / Delivery man.
Tool Use 288 / 197 Q: How do I turn on the stove?
A: Turn the knob on the front of the stove clockwise until you feel a click. / By clock in the switch button on the stove. / Rotate the button below. / Turn the switch to the left.
Other
Resources
366 / 258 Q: Is there anyone nearby that I can ask for directions?
A: Yes. / Yes, there is a person sitting in the front. / Yes, there is a security guard ahead. / Yes, there is front-desk security.
Manual Extraction. We first manually watch the video clips and extract the vision-assistive questions from the camera wearers, the timestamps of the questions are also recorded. It is worth noting that these in-video questions are answered by the blind users’ sighted partners or others in the videos. Thus, to prevent answer leakage and ensure real-time QA, models can only access visual content up to the question timestamps to answer a specific question. We obtain a total of 500 questions during this process since qualified questions are sparse in the videos.
AI-Generation. To enrich the annotations, we consider generating questions by prompting GPT-4o [29] to act as blind questioners, followed by rigorous verification with real blind individuals. Specifically, we first categorize the assistance scenarios into six types based on our observation of the video content and the extracted questions: Information Reading, Safety Warning, Navigation, Social Communication, Tool Use, and Other Resources. Typical question of each group is listed in Tab. 1. To limit generative hallucinations and ensure a diverse set of examples for each assistance type, QA generation is conducted categorically, focusing on one assistance type at a time. Moreover, the generation and verification are conducted alternately to ensure high generation quality.
Specifically, in our prompting, the generated QA pairs are required to adhere to the following criteria to ensure that they are practical, context-aware, and closely aligned with the needs of the visually impaired individuals: 1) First-Person Perspective: All questions should be framed from the perspective of the visually impaired individuals (camera wearers). 2) Practicality: The questions must be practical and reflect the blind’s true needs in the given situation. 3) Video-Level Questions: Questions should reflect the temporal aspect of the video, requiring multiple frames to derive the answer and emphasizing dynamic characteristics. 4) Real-Time Context: Questions must pertain to the real-time context of the video and exclude any external information beyond the visible content before and at the question time stamp. Other details are presented in Appendix A.
3.1.3 Manual Verification
Our verification is done in three stages involving both sighted and visually impaired people. Also, the first three authors are engaged into all stages to control the quality. Concretely, in the first stage, the authors are responsible to check the QA quality every after generating a limited number of questions for one type of assistance. Related issues are recorded and reflected in the prompts for subsequent generations, thus to maximally ensure the generation quality. In such an alternate way, we generate 16,217 questions.
In the second stage, we invite 63 volunteers (sighted people) to thoroughly review the questions. In this process, we remove lots of questions that are 1) redundant in meaning; 2) not easy to answer (e.g., “How do I use my phone?”); and 3) meaningless (e.g., “How do I use my white cane?”). We also find and delete a moderate number of questions that are 1) not related to vision (e.g., “How heavy is this box?”) and 2) not from first-person perspective (e.g., “what is the man (camera wearer) doing?”) . Note that all ambiguous questions are further cross-validated by the three authors to either remove or edit them. What remained are 4,737 questions after this stage, in which 82.8% are edited by humans. Additionally, the 500 extracted questions also undergo checking for translation errors, resulting in a total of 5,237 valid questions after this stage.
In the third stage, we sample 30% data (1.5K) corrected questions and invite 111 visually impaired volunteers (with diverse backgrounds) to review them to ensure their relevance to the blind users’ true needs for visual assistance (see Appendix A). The result observations are then reflected throughout the whole dataset by the authors. Specifically, the blind volunteers are requested to score the confidence (from the lowest 0 to the highest 5) of a question that it may be asked under a specific visual scenario. Each question will be scored by at least 10 volunteers and maximally 33. Additionally, the bind volunteers are encouraged to pose new questions optionally. With the study result, we further remove the questions with an average confidence score of lower than 1 and add a limited number of new questions (about 300) annotated by the blind volunteers. Other details are presented in Appendix A. Eventually, we obtain 4,927 valid questions. Their distributions are listed in Tab. 1.
3.1.4 Answer Annotation
We invite 21 university students to watch the videos and answer the questions. To enhance the persuasiveness of the standard answers and reduce subjective evaluation, each question is assigned to three or four annotators for answering. Only one may see the generated answer (if available) and must verify and edit it as needed. The others are required to directly answer the question for assistance purpose. If there is no generated answer, e.g., for the manually extracted or annotated question, the annotators need to answer the question themselves. We resolve contradictory answers through consensus between the authors and the majority of annotators. Redundant answers are avoided by merging similar responses. Ultimately, each question has 1 to 4 valid answers (3 on average). Additionally, about 67.7% of the generated answers are edited by human annotators, and most unedited answers are deterministic, such as yes/no and numbers. Examples of the reference answers are shown in Tab. 1.
Figure 3: Examples of EgoBlind.
3.2 Statistic Analysis
Data Split. The annotation process takes about 10 months. We finalize 4,927 questions across 1,210 videos (two examples are shown in Fig. 3). We split the dataset into training and test sets at a 7:3 ratio while ensuring that video clips from the same source do not appear in both. As shown in Tab. 2, the training set contains 906 videos and 3,438 questions, while the test set includes 304 videos and 1,489 questions. Additionally, we employ Gemini 2.0 to probe if there is a significant discrepancy of model behaviors between the test set and the whole dataset. We find that the prediction accuracy differences are quite small, ranging from 0.3% to 0.8% across different question types. This suggests that the test set is sufficient to serve as a reliable testbed for evaluating model behaviors.
Table 2: Statistics of the EgoBlind datasets
Videos Questions
Train Test All Train Test All
906 304 1,210 3,438 1,489 4,927
(a) Distribution of question length.
(b) Distribution of answer length.
(c) Top-10 frequent questions.
(d) Top-10 frequent answers*.
(e) Word cloud of questions.
(f) Word cloud of answers*.
(g) Distribution of video durations.
Figure 4: Statistic analysis of EgoBlind. *: We omit 1,123 (22.8%) “yes/no” and 541 (11.0%) “do not know” answers in (d) and (f) for better presentation. (Please zoom in for better view.)
Questions. Fig. 4(a) shows that more than 80% of the questions are short in 10 words, with an average length of 7.9. The relatively shorter questions could be attributed to the use of more spoken language in practice. The word cloud in Fig. 4(e) shows that the questions are from blind users’ first-person perspective (“I”, “my”, “me”) and feature non-verbal concepts, such as “this”, “there”, “that”, “it” and “now”, which often demand on engaging in the scenarios, capturing the questioner’s hand pose and eye gaze, and reasoning the spatio-temporal contexts for understanding the questions. Fig. 4(c) shows that blind people are often interested in locating something (“where is”) and checking for potential safety issues (“are there any obstacles”).
Answers. Fig. 4(b) shows that the answers are relatively longer with an average of 5.7 words. Near half of the answers have more than 3 words, with 19.4% over 10 words. A detailed study shows that the answers for tool use are significantly longer and more complex than other assistive questions; the average answer length is 12.1 for tool use while 4 to 6 for others. The most frequent answers are “yes” (13.7%), followed by “I do not know” (11.0%) and “no” (9.7%). The answer of “I do not know” is due to poor visual quality or the answer not framed by the blind users. Additionally, it may attribute to our real-time QA task setting, i.e., the answer is not visible up to the question moment. By omitting these answers, we analyze the remaining frequent answers and answer word clouds in Fig. 4(d) and Fig. 4(f), respectively. We find that the answers are mostly about directions, navigation and locations. It is worth mentioning that the locations and directions are often relative to the blind camera wearers (questioners).
Videos. Fig. 4(g) shows that the vast majority of the videos (84.3%) are under 1 minute, with 15.7% exceeding 1 minute and 5.9% running longer than 2 minutes. The mean video duration is 35.8 seconds, reflecting the average duration each time a visually-impaired person recording their daily moment. Additionally, we conduct a comparison between the egocentric video content captured by sighted people (e.g., Ego4D [13] videos) and those are blind. We find that the differences are from the following major aspects: 1) Composition and Focus: Blind individuals often produce footage where subjects are off-center or out of focus (refer to Fig. 1 and Fig. 3). 2) Camera Orientation and Stability: Blind individuals hard to maintain consistent camera orientation, potentially leading to tilted or unstable footage. 3) Environmental Awareness: Blind individuals inadvertently capture obstructed views or poorly lit scenes. Nevertheless, the video resolution is high, possibly because our videos are crawled from video sharing platforms.
3.3 Dataset Comparison
The EgoBlind dataset, as a first-person perspective video question-answering dataset, has distinct uniqueness. Unlike most first-person perspective QA datasets (datasets listed in the first block of Tab. 3), which focus on assessing general-purpose visual understanding abilities, the core task of EgoBlind centers around assisting blind individuals, emphasizing the real-time and practical nature of tasks. Based on the actual needs of blind people, we categorize the questions to evaluate the model’s ability in different assistance aspects. Compared to other assist-oriented datasets (listed in the middle block of Tab. 3), EgoBlind covers the wide aspects of visual assistance for the blind people, versus episodic memory in QAEgo4D [3] or tool use in AssistQ [36] or scene-text QA in EgoTextVQA [48].
Compared to the blind-related QA datasets (such as the well-known VizWiz [14] image dataset and the recent video QA dataset VIEW-QA), EgoBlind advances VizWiz by enabling egocentric live QA with more visual information about real-world dynamics and episodic memory support. Although the videos in the VIEW-QA [33] dataset offer 360-degree visual information, the videos and questions are simulated by people with normal eyesight mimicking the perspective of blind individuals, thus lacking sufficient practical characteristics for the blind community, especially for those with long blind age. In contrast, the videos in EgoBlind are entirely filmed by blind individuals themselves, authentically reflecting their real-life scenarios. Moreover, the questions in EgoBlind are posed or verified by blind individuals of different blind ages, further ensuring the dataset’s practicality and relevance.
Table 3: Dataset Comparison. BlindV: Video captured by blind people. BlindQ: Question posed or verified by blind people. RTime: Answer are based on the video content up to the question timestamp. MA: Multiple reference answers for each question. QC: Question Category. OE/MC: Open-ended/Multi-choice.
Datasets BlindV BlindQ RTime MA QC Task
General-purpose EgoVQA
EgoVQA [10] OE
EgoTaskQA [17] OE
EgoThink [7] OE
VidEgoThink [6] OE
EgoSchema [28] MC
EgoMemoria [42] MC
Assistance EgoVQA
QAEgo4D [3] OE
AssistQ [36] MC
EgoTextVQA [48] OE
Blind VQA
VizWiz [14] - OE
VIEW-QA [33] OE
EgoBlind (Ours) OE
4 Experiment
4.1 Experimental Setups
Evaluation. We used GPT-4o mini as a validation tool to assess the model performances on EgoBlind test set. It evaluates the semantic similarity between the models’ predictions and the ground truth (reference answers). Building on previous work, we established two evaluation metrics for the models’ predictions: accuracy (0-100%, the percentage of ‘yes’ answers evaluated) and score (0-5, degree of similarity to the matching results). Specially, for each test question, a correct prediction is identified if it achieves yes-response with any one of the reference answers. It is worth mentioning that the evaluation prompts are manually finetuned to reach a maximal score agreement (0.88) between human and AI reviewers (Details are introduced in Sec. B.1). The final evaluation prompts are attached in Tab. 9.
Table 4: Performance (average accuracy and score) of sampling uniformly across the video content up to the question timestamp (video input) vs. sampling 1 frame at the question timestamp (image input). Sampling multiple frames bring better performances.
Methods Video Input Image Input
Video-LLaVA [22] 38.5 / 2.1 36.7 / 2.1
LLaMA-VID [21] 39.2 / 2.2 35.6 / 2.1
LLaVA-OV [19] 55.5 / 2.9 54.2 / 2.8
InternVL2-8B [5] 56.5 / 3.0 54.7 / 2.9
Gemini2.0[34] 56.6 / 3.0 52.4 / 2.7
Model Setup To benchmark the challenges embodied by EgoBlind, we comprehensively analyze 15 contemporary MLLMs, including 12 open-source models and 3 closed-source models (via APIs). The choice of the open-source models are based on the following criteria: 1) Achieve SOTA results on common video QA benchmarks such as ActivityNet-QA [43], NExT-QA [37], VideoChatGPT-Bench [27] and Video-MME [11]. Corresponding models are Qwen2-VL [35], LLaVA-OV [19], CogVLM2-Video [15], Video-LLaMA3 [44], InternVL-2 [5]; 2) Achieve SOTA results on recent general-purpose egocentric VideoQA benchmarks such as EgoSchema [28], VidEgoThink [6] and EgoMemoria [42]. Corresponding models are LLaVA-NeXT-Video [19], MiniCPM-V 2.6 [41] and Qwen2-VL[35]; 3) Achieve SOTA on the image blind QA dataset VizWiz [14], such as Video-LLaVA [27], LLaMA-VID[21] and VILA 1.5 [23]. We additional include ShareGPT4Video [4] as it pioneers in using Ego4D videos for pretraining.
Noteworthy, for each question, we uniformly sample the video content up to the question timestamp for predicting their answers, thus to meet the live QA task setting. Additionally, we have experimented an alternative video sampling strategy by using the frame at the question timestamp, and find that it jeopardizes the performances (refer to Tab. 4). Also, we design customized prompts (refer to Appendix Tab. 7) tailored for answering blind users’ questions for better performance (see analysis in Sec. B.2). Other details follow the official model protocols.
Table 5: Model behaviors on EgoBlind. No single model wins across all question types. Models that are superior on general-purpose egocentric VQA or image blind QA are not the best-performing. The closed-source models perform similarly to the top-performing open-source ones. All results significantly lag behind human performance.
Methods LLM Res. #F Tool Use Information Navigation Safety Communication Resource Average
Human - - - 70.4 / 3.4 87.0 / 4.3 83.1 / 3.9 91.9 / 4.5 94.7 / 4.7 96.6 / 4.6 87.4 / 4.2
Open-source Models
ShareGPT4Video [4] LLaMA3-8B - 2fps 32.2 / 2.3 32.5 / 1.8 14.8 / 1.4 40.7 / 2.2 31.6 / 1.9 30.8 / 1.8 33.3 / 2.0
CogVLM2-Video [15] LLaMA3-8B 2242 24 35.6 / 2.3 42.7 / 2.3 15.7 / 1.4 50.0 / 2.6 60.1 / 3.2 33.3 / 2.0 40.0 / 2.3
Video-LLaMA3 [44] Qwen2.5-7B ori 1fps 45.6 / 2.5 54.0 / 2.7 34.8 / 2.0 48.4 / 2.6 55.3 / 2.8 56.4 / 2.9 49.6 / 2.6
InternVL2-26B [5] InternLM2-Chat-20B 4482 8 58.9 / 3.1 55.5 / 2.9 51.4 / 2.8 54.1 /2.9 47.4 / 2.6 59.8 / 3.1 54.9 / 2.9
LLaVA-OV [19] Qwen2-7B 3842 32 57.8 / 3.2 58.5 / 3.0 30.1 / 2.0 64.2 / 3.3 63.2 / 3.2 51.3 / 2.8 55.5 / 2.9
InternVL2-8B [5] InternLM2_5-7B-Chat 4482 8 61.1 / 3.3 57.2 / 3.0 48.1 / 2.7 59.3 / 3.1 47.4 / 2.5 58.1 / 3.1 56.5 / 3.0
MiniCPM-V 2.6 [41] Qwen2-7B 3842 1fps 48.9 / 2.6 47.6 / 2.5 43.3 / 2.4 32.0 / 1.9 47.4 / 2.6 45.3 / 2.4 43.0 / 2.3
Qwen2-VL [35] Qwen2-7B ori 1fps 53.3 / 3.0 50.0 / 2.6 38.6 / 2.3 35.8 / 2.0 47.4 / 2.5 50.4 / 2.7 45.1 / 2.4
LLaVA-NeXT-Video [19] Qwen2-7B 3842 1fps 41.1 / 2.4 54.8 / 2.8 35.2 / 2.1 62.0 / 3.2 52.6 / 2.9 54.7 / 2.8 52.9 / 2.8
Video-LLaVA [27] Vicuna-7B 2242 8 27.8 / 1.9 40.7 / 2.2 18.1 / 1.3 48.9 / 2.5 52.6 / 2.8 33.3 / 2.0 38.5 / 2.1
LLaMA-VID [21] Vicuna-7B 2242 1fps 35.6 / 2.2 39.8 / 2.2 30.1 / 2.0 47.8 / 2.5 50.0 / 2.7 36.8 / 2.1 39.2 / 2.2
VILA1.5 [23] LLaMA3-8B 3362 32 41.1 / 2.4 52.4 / 2.7 26.2 / 1.7 56.0 / 2.8 60.5 / 3.2 40.2 / 2.4 48.2 / 2.6
Closed-source Models
Gemini 1.5 Flash - ori - 61.1 / 3.3 57.9 / 3.0 38.1 / 2.3 57.3 / 2.9 60.5 / 3.0 53.0 / 2.9 54.8 / 2.9
GPT-4o - ori 1fps 60.0 / 3.2 55.8 / 2.9 47.6 / 2.6 59.9 / 3.0 48.7 / 2.6 59.8 / 3.1 56.0 / 2.9
Gemini 2.0 - ori - 48.3 / 2.9 57.0 / 3.0 51.1 / 2.8 47.7 / 2.7 63.2 / 3.2 65.2 / 3.4 56.6 / 3.0
4.2 Benchmarking Analysis
4.2.1 Overall Observations
Table 5 presents the performances of different models across different question types. We outline the following main observations below:
[…]
  • The closed-source models perform similarly to the top-performing open-source ones such as InternVL2 [5] and LLaVA-OV [19].
  • The models that are superior at general-purpose ego-VQA and image blind-QA are not the best-performing.
  • No single model wins across all question types.
  • Answering “Navigation” questions is the most challenging task for almost all models, followed by “Tool Use”.
  • Better LLMs always bring but are not clear-cut for better performance (Qwen2 vs. LLaMA3 vs. Vicuna).
  • Large variant of the same models (e.g., InternVL2-26B vs. 8B) may not necessarily be the stronger one.
4.2.2 Specific Challenges
Figure 5: MLLMs face challenges in understanding user intention. The question is posed at the last frame (same as below).
User Intention. We found that, even when the prompt emphasizes that the model should serve blind people, the models still fail to fully understand the context and intent of the question in real scenarios, leading to inaccuracies in its answers. For a typical example in Fig. 5, where a blind person asks, “Is there a road ahead?” although a road is visible in the image, the subject’s path is actually blocked by a number of obstacles, such as yellow-black roadblocks, lamp posts, and large trees. A reasonable response should be “No,” with further recommendations for an alternative path, as suggested by all the ground-truth answers. However, the model generally tends to answer “Yes”, without considering the impact of these obstacles for navigation. For improvements, it is crucial to spatio-temporally engage into the users’ real-time activities.
Figure 6: MLLMs face challenges in reasoning spatial orientation.
Real-time Spatial Orientation. MLLMs exhibit significant shortcomings in understanding spatial orientations relative to the subjects (blind people) when they are moving. A typical example is shown in Fig. 6. Specifically, when the subject moves, the models fail to effectively update their real-time spatial orientations relative to the subject (e.g., “right side”), and favor static and prominent objects (e.g., “oil”) while overlooking the actual needs from the subjecs’ perspectives (e.g., “tomato”). A similar case is also found in Fig. 8, where all models (except for Gemini 2.0) fail to assist the user in localizing the escalator. This posits to a limitation towards egocentric embodied AI assistance. For improvements, the models should be sensitive to the subtle visual difference in continuous video frames.
Temporal Context. MLLMs often fail to reason the temporal contexts to precisely understand and answer the question. For example in Fig. 7, the blind user asks a simple question “where is my companion?”. Only Gemini gets close responses, possibly because of its raw video inputs (continuous frames and audio). Most models fail to correctly answer the question. It seems that the models do not know who is the companion of the blind user, and hallucinate their answers that mismatch the visual fact (e.g., “on the boat”, “on the left”). For human, it is easy to know that the “man in grey” is the user’s companion even by a simple glance over the video content. Similarly, Fig. 8 demonstrates that most models fail to recall the temporal moment (the 1st frame) that shows the location of the escalator. This discrepancy reveals a model limitation in reasoning the temporal contexts. Long-memory techniques could help address this limitation.
Figure 7: MLLMs fail to reason the temporal contexts.
Figure 8: MLLMs fail to reason the spatio-temporal contexts.
Identifying Obstacle. MLLMs’ performance in “safety warning” questions remains inadequate. They cannot serve as reliable safety tools to provide users with accurate answers. For example, in a typical scenario where the subject is walking along the road and asks “Are there any obstacles in my path?” the answers generated by the models often show significant deviations from the real situation (Ground Truth, GT), exhibiting a mismatch with the question asked. Even though there may be potential obstacles in the scene (such as pedestrians, vehicles, or other barriers), the models fail to accurately identify and alert the user, instead providing irrelevant or vague responses. This discrepancy not only affects the accuracy of the answers but also reflects a mismatch between the style of the answers and the user’s actual needs. Despite having strong scene detection abilities, the models still fall short in understanding the user’s intent within context, identifying potential risks, and providing precise warnings. For improvement, we hope our training data can serve as a valuable resource.
Figure 9: MLLMs face challenges in identifying obstacle.
Figure 10: MLLMs tend to be sycophant.
Sycophancy. MLLMs exhibit significant shortcomings when the blind users ask non-existent object due to blindness. For example, in a typical scenario where a blind person, sitting in a car with no other passengers, asks “Who is sitting next to me?”, only human could correctly understand the intent and provide a negative answer (e.g., “No one is sitting next to you”). In contrast, the models often deviate from the actual visual situation or the question. For instance, some models hallucinate that there are other passengers in the car, while others identify the driver as sitting next to the user but fail to give a negative answer. This sycophantic phenomenon highlights the limitations of current MLLMs when dealing with unconventional questions from blind users. Due to their visual impairment, blind individuals often ask questions that may seem “strange” to sighted individuals but are entirely reasonable from their perspective. For alleviation, synthesized data for tailored training could be helpful.
5 Conclusion
In this paper, we present the construction of a VideoQA dataset EgoBlind to reflect on the challenges towards egocentric visual assistance for the blind people. EgoBlind is the first of its kind in that the videos are token by the visually-impaired individuals from a first-person perspective, with the questions posed and verified by blind users to ensure close alignment with their true assistant needs. We further comprehensively benchmark the challenge with 15 prominent multimodal LLMs and unveil their significant limitations in various aspects. We conduct thorough analysis and share many key insights for future advancements in this direction. By formulating and bringing the challenge to the vision-language community, our primary goal is to push the MLLM research towards live egocentric visual assistance for the overwhelming number of visually impaired people around the world.
Importantly, we have ensured that the dataset collection and user studies adhere to IRB standards. Informed consent was obtained from all the content creators, who agreed to their content being used for non-commercial research purposes. Access to the dataset will be granted via authorized signatories, and video sources will be cited for distribution.
Appendix
Appendix A EgoBlind Dataset Creation
A.1 AI Generation
The process of calling the GPT-4o API to generate QA pairs is mainly divided into three parts. Video Decoding and Frame Extraction. We first perform video abstraction on the edited video clip by saving one frame every 10 frames. The process of generating QAs from video involves several steps: The video is decoded into individual frames. For efficiency, every 10th frame is extracted (saved), which reduces the amount of data while retaining key visual information. Uniform Sampling. After initial frame extraction, further uniform sampling is applied. The sampling interval varies based on the video length to balance resource usage and data redundancy. For instance: A 3-minute video might be sampled every 18 frames. A 30-second video might be sampled every 6 frames. This adaptive approach ensures that shorter videos have more frequent samples, capturing essential details, while longer videos maintain a broader overview. GPT-4o Model Input The uniformly sampled frames are input into the GPT-4o model. Each frame is accompanied by its timestamp and a refined prompt (as shown in LABEL:tab:gen_qa), instructing the model to generate relevant, real-time question-answer pairs based on the frame content. Batch Processing for Diversity. To enhance data diversity, the generation process is divided into batches by question categories. This means that each batch contains only one type of question, which can help in organizing and enriching the dataset, ensuring a wide coverage of different topics and improving overall quality. This structured approach ensures that the generated QA pairs are both relevant and diverse, effectively utilizing video content to create a comprehensive and organized dataset. The specific prompts for each category are attached in LABEL:tab:gen_qa.
A.2 Blind User Study
We categorize manually corrected QA pairs (translated into Chinese) based on video scenes and invite 111 blind individuals to evaluate the blind-relevance of the questions across different scenarios. The evaluation scale ranges from 0 to 5, with higher scores indicating preferable questions and lower scores reflecting unwanted questions (as illustrated in Fig. 13, a real example is shown in Fig. 20). Additionally, blind individuals are encouraged to provide further questions and insights to better meet the needs of individuals with visual impairments.
Figure 11: Distribution of blind participants.
The cohort study (Fig. 11) exhibits substantial demographic heterogeneity, encompassing participants ranging from teenagers to older adults (Fig. 12). Male participants account for a dominant proportion of 73.3%, with young adults aged 18-30 constituting the largest age cohort. Notably, for most respondents, the duration of visual impairment corresponded with their biological age, suggesting congenital origin - a finding consistent with the observation that over 80% presented Grade 1 Blindness (total visual acuity $\leq$0.02 or visual field <5absent5<5< 5°). The remaining participants acquire visual impairment postnatally, though all reported living with the condition for multiple years.
The occupational distribution reveals a pronounced concentration in specialized massage services (55.6%), reflecting a predominant vocational pattern among the visually impaired population of China. Students form the secondary occupational group (18.9%), followed by unemployed individuals (10.0%) and other professions (15.5%). Geographically, the sample represents 46 urban centers across China’s hierarchical city-tier system, ranging from first-tier megacities (Beijing, Guangzhou, Shenzhen) to emerging technology hubs (Hangzhou, Chengdu, Changsha) and smaller urban cities (Neijiang, Zhongshan, Langfang).
This demographic and geographic diversity ensures comprehensive coverage of life scenes across varying urban contexts, from high-density metropolitan transit networks to community-level infrastructure systems.
Figure 12: Distribution of user ages and their blind ages.
Figure 13: The scoring example
We found that when blind people visit supermarkets, stores, and shopping malls (as shown in Fig. 14); their primary concerns are navigation and positioning issues related to location. Specifically, they want to know where they are and how to get to their desired destination(“How to get to XXX from here?”) . Their second main concern is obtaining product information, such as price (“Where can I find the price of this product?”), style, shelf life, etc. Additionally, some blind individuals are also interested in understanding the range of products sold at specific stores and the layout of the mall, including questions like “What does this store sell?” and “Which products are displayed on the counter?” At the same time, we found that blind individuals tend to be less concerned about the presence and location of people around them. Questions related to the number of people nearby (“How many people are in front of me?”) and their specific whereabouts (“Is the cashier directly in front of me?”) received a significantly higher proportion of low scores compared to high scores. Additionally, they showed minimal interest in information about road obstacles. The proportion of low scores for questions such as ”What are the obstacles on the road?” and ”How many floors are there?” was considerably higher than the proportion of high scores.
Figure 14: MOS score distribution of blind people in shopping scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 15: MOS score distribution of blind people using transportation. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 16: MOS score distribution of blind people navigating outdoor environments. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 17: MOS score distribution of blind people in shopping scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
Figure 18: MOS score distribution of blind people in cooking and doing craft scenarios. Each row of the ordinate represents a question, and the stacked parts of each color represent the distribution proportions of different scores.
When blind people use transportation such as cars, buses, and other vehicles (as shown in Fig. 15), over 60% of the respondents assigned full scores to questions like ”Is this my taxi?” and ”Which bus is coming?” while waiting for a taxi. This indicates that their primary concern is the current location of the vehicle. Meanwhile, respondents uniformly gave low scores to the question ”Is the car being moved?” suggesting that blind individuals are capable of perceiving vehicle movement. Furthermore, approximately 80% of respondents gave low scores to questions such as ”What does the bus stop advertisement say?” and ”What brand is the car?” indicating that blind people often show little interest in irrelevant information surrounding their waiting position.
In outdoor navigation scenarios, such as finding destinations, crossing roads, or walking, blind individuals prioritize safety above all else (as shown in Fig. 16). For instance, the question “Is it safe to cross the road?” receives the highest score of 5 from all respondents. Similarly, over 80% of respondents rate questions related to safety during movement, such as “How can I safely get through this entrance?” and “Is it safe to go straight?”, with scores of 4 or higher. Blind individuals also place significant emphasis on directional and positional information. For questions about accessing public facilities, such as “How to reach the nearest sidewalk, zebra crossing, or blind road?”, over 70% of respondents gave scores of 4 or higher. Direction-related questions like “Which direction should I take to reach the intersection?” and “Is the supermarket on my left?” received scores above 3 from more than 60% of respondents. However, respondents show less concern for questions about the characteristics of the surrounding environment, such as “Please describe the environment around me” and “Is traffic busy in this area?”. Similarly, questions about weather conditions (“Is it sunny now?”) or information about nearby pedestrians (“Who is the person standing in front?”, “Are there many people around us gathered?”, “Is there a pedestrian coming towards me?”) received scores below 3 from more than 50% of respondents, indicating a general lack of interest. Interestingly, questions about obstacles did not receive as high a score as might be intuitively expected. For example, fewer than 50% of respondents positively rated questions like “Can you describe the obstacles ahead?” and “To avoid obstacles, which direction should I go?”, while a large proportion gave neutral scores (3). This may be due to differences in the proficiency of using white canes among blind individuals, which influences their reliance on such information. Similarly, questions such as “Am I walking on a blind road” also exhibited a polarized response, with respondents either assigning very low scores (0 or 1) or very high scores (4 or 5). This polarization suggests that reliance on certain types of information may vary significantly depending on individual preferences and skills.
We also observed characteristics of blind individuals’ activities in familiar environments that contradict common assumptions. As shown in Fig. 15, when blind individuals are at home, we found that the proportion of questions receiving low scores (0, 1, or 2) exceeded 60% across almost all queries. Respondents explained that although they cannot see, their brains form a mental “ma” of the environment based on prior familiarity. This mental representation allows them to navigate and plan their movements seamlessly within the environment. Unless there are significant changes in the environment, they typically do not encounter any issues.
To validate the rationality of question design in hands-on activity scenarios for visually impaired individuals, we conducted a study where participants rated questions in scenarios such as cooking and crafting (As shown in Fig. 18). Taking these two scenarios as examples, we found that, despite the differences in context, respondents consistently gave high scores (4 and 5) to core questions that directly influence task completion. For instance, “How much time is left?” (Cooking) and “Please describe the items on the shelves” (Doing crafts). At the same time, due to the differences in activities, visually impaired individuals tended to give a higher proportion of low scores (0 or 1) to cooking-related questions compared to crafting. This can be attributed to the fact that cooking typically occurs in familiar environments where blind individuals rely on their spatial memory and experience. In contrast, in the crafting scenario, respondents showed a neutral or low level of interest (0–3) in questions about basic item information such as appearance or shape, exemplified by questions like “Is it a cup?” and “What does this cup look like?”.
Appendix B Experimental
B.1 Evaluation
To enhance the reliability of the AI evaluation tool (i.e., GPT-4o mini), we sample 30% of the test data to obtain inference results of Gemini 2.0. We invite human volunteers to score the models’ inference results on this subset. Based on the results of human scoring, we refine the prompts to guide the machine’s evaluation results closer to the human scoring outcomes. Ultimately, we calculate the Cohen Kappa coefficient [18] (for inter-rater reliability) between human and GPT scoring results, which is found to be 0.73, indicating a high degree of consistency between the two. Through heatmap analysis (shown in Fig. 19), we observe that samples where human and machine ratings are consistent accounted for 88% of the total test data. For samples where ratings are inconsistent, we find that most of they are due to an information bias – machine cannot see the videos for answer scoring.
Figure 19: Consistency between human and machine scoring.
B.2 Other Investigations
We conduct targeted experiments to compare the effects with and without specific prompts for the blind. Tab. 6 shows that designing special prompts for blind users can effectively improve model performances, versus naively prompting the models to answer the questions.
Table 6: Effectiveness of our prompt instructions customized for blind people. If without prompting the models that they are answering questions for assisting blind people, remarkable performance drops are observed almost on all question types.
Method Input Tool Use Information Safety Communication Navigation Resource Average
Gemini2.0 Blind-specific 48.3 / 2.9 57.0 / 3.0 51.1 / 2.8 47.7 / 2.7 63.2 / 3.2 65.2 / 3.4 56.6 / 3.0
[…]
InternVL2-8B Blind-specific 61.1 / 3.3 57.2 / 3.0 48.1 / 2.7 59.3 / 3.1 47.4 / 2.5 58.1 / 3.1 56.5 / 3.0
[…]
Figure 20: Exampe of questionnaire for blind study. Questions are translated into Chinese.
Model General Prompts
Open Source MLLMs I will provide you with a video each time and one question; These questions are all questions raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question which was posed in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. The question is: {question}
GPT-4o I want you to act as a blind person. I need help identifying blind people. I will provide you with several pictures each time and one question; These questions are all raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {’question’:’Is there a wet floor caution sign near the door?’, ’question_id’: ’v_87_1_2’, ’prediction’:’yes.’, ’type’: ’toolrmation reading’}
Gemini I want you to act as a blind person. I need help identifying blind people. I will provide you with a video each time and one question; These questions are all questions raised by the blind person in the video from his own first-person perspective in the current scene. Your task is to answer the blind person’s question which was posed in the last frame of the video. The answer needs to be based on the content of the picture and the objective characteristics that the blind person cannot see. If the question cannot be answered, you can say I don’t know. Do not include Chinese characters in your response. Please generate the response with keys ’question’, ’question_id’, ’prediction’, and ’type’. The question is: {question},Question ID: {question_id},Type: {question_type}. The content of the key ’question’ is the question you received. The content of the key ’prediction’ is the answer to the corresponding question you generated. The key ’question_type’ represents which category the question type you received. Your response should look like this example: {”question”:”Is there a wet floor caution sign near the door?”, ”question_id”: ”v_87_1_2”, ”prediction”:”yes.”, ”type”: ”information reading”}
Table 7: Prompts for models to perform question answering on EgoBlind.
Table 8: Prompts for generating different types of QAs in EgoBlind.
Question Category General Prompts
Social Communication I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ‘communication and interaction’.Below I will give a detailed explanation of ‘communication and interaction’. Communication and Interaction:Describe the interaction between blind people and surrounding people/companions, and observe other people‘s status information for multi-person collaborative activities. For example:Did my companion help me carry my luggage onto the conveyor belt? /Has my guide dog entered the elevator safely? /Who is talking to me? Here’s how you can accomplish the task:1.use your imagination and only ask about ‘communication and interaction’questions;2.Please ask questions about each picture;3.The questions you ask need to be real-time and do not contain any off-site information;4.The questions you ask need to be answered through picture content;5.If you cannot answer based on the picture, you can answer ‘I don’t know’;6.Ask as many open-ended questions as possible, not just ’yes-no’ questions;7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is my companion doing? 8.‘communication and interaction’questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation).9.Please don’t ask bad examples like ‘What is my companion saying to me ?’/‘Is there an announcement about the next station or any other important information?’ Please generate the response with ’timestamp’, ’question’, ’answer’ and ’questoin_type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.The timestamp value can be found here:’f’Timestamp list: time.
Information Reading I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of information reading which called ’information reading’. Below I will give a detailed explanation of ’information reading’. ”Information Reading:Used to obtain some kind of information you want to know. For example:What is the menu about?/What is my mobile phone model?/I just pressed the elevator button. Could you please tell me which floor I just pressed?” Here’s how you can accomplish the task: 1.use your imagination and only ask about ’information reading’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’information reading ’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys, ’timestamp’, ’question’,’answer’and ’questoin-type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’information reading’,your response should look like this example:’timestamp’:’3000’,’question’:’What is this?’,’answer’:’It is a bag.’,’question_type’:’information reading’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Navigation I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. You need to focus on asking questions about the types which called ’local navigation and positioning’.” ”Below I will give a detailed explanation of ’local navigation and positioning’.” ”Local Navigation and Positioning:Find the correct direction/location/navigation information needed by blind people, focusing on simulating indoor navigation and outdoor local navigation functions. For example:where the toilet is? /Am I facing the gate now? /how to get to the nearest stairway from here?” ”Here’s how you can accomplish the task: 1.use your imagination and only ask about ’local navigation and positioning’ questions; 2.please ask questions about each picture; 3.The ’local navigation and positioning’ questions you ask need to be real-time and do not contain any off-site information; 4.The ’local navigation and positioning’ questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended ’local navigation and positioning’ questions as possible, not just ’yes-no’ questions. 7.Some of the ’local navigation and positioning’ questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’local navigation and positioning’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys, ’timestamp’, ’question’,’answer’and ’questoin-type’. The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’local navigation and positioning’,your response should look like this example:’timestamp’:’3000’,’question’:’Where is the crosswalk?’,’answer’:’It is in front of you.’,’question_type’:’local navigation and positioning’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Safety Warnings I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided.You need to focus on asking questions about the types of tools you use which called ’safety warnings’. Below I will give a detailed explanation of ’safety warnings’. Safety Warnings:Help blind people determine whether their environment has physical safety hazards/human factor safety risks.For example: Is there any traffic passing by when crossing the road?/What is the direction of the escalator?/Are there any roadblocks during the walk?/Are there any malicious people nearby? Here’s how you can accomplish the task: 1.use your imagination and only ask about ’safety warnings’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’ 8.’safety warnings’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’, ’question’, ’answer’ and ’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’safety warnings’your response should look like this example:’timestamp’:’3000’,’question’:’Is it safe to cross the road now?’,’answer’:’yes.’,’question_type’:’safety warnings’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Tool Use I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ’tool use’.Below I will give a detailed explanation of ’tool use ’. tool use:Describe how to use a tool or operate a program. For example: How to use microwave oven?/How to use induction cooker and dishwasher? /How to operate the gate? Here’s how you can accomplish the task:1.use your imagination and only ask about ’tool use’ questions;2.please ask questions about each picture;3.The questions you ask need to be real-time and do not contain any off-site information;4.The questions you ask need to be answered through picture content;5.If you cannot answer based on the picture, you can answer ’I don’t know’;6.Ask as many open-ended questions as possible, not just ’yes-no’ questions.7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’How should I do?’8.’tool use’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’,’question’,’answer’and’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’tool use’your response should look like this example:’timestamp’:’3000’,’question’:’How to operate the gate?’,’answer’:’go straight ahead.’,’question_type’:’tool use’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Other Resource I want you to act as a blind person. I need help identifying blind people.I will provide you with several pictures each time and the pictures have two characteristics:1. pictures have temporal correlation between them and the order of the pictures represents the order of time. 2.The pictures are all taken from your first point of view, which is equivalent to the picture recorded by the camera hanging on your chest. Your task is to imagine what questions you would ask to meet your needs if you were in the situation of the picture provided. You need to focus on asking questions about the types of tools you use which called ’resource acquisition and others’. Below I will give a detailed explanation of ’resource acquisition and others’. Resource Acquisition and Others:Looking for nearby services facilities or activities that blind people would like. For example: Is there barrier-free blind walk/manual service/a certain brand store ? ”Here’s how you can accomplish the task: 1.use your imagination and only ask about ’resource acquisition and others’ questions; 2.please ask questions about each picture; 3.The questions you ask need to be real-time and do not contain any off-site information; 4.The questions you ask need to be answered through picture content; 5.If you cannot answer based on the picture, you can answer ’I don’t know’; 6.Ask as many open-ended questions as possible, not just ’yes-no’ questions. 7.Some of the questions generated need to reflect the characteristics of the video, that is, you need to watch several consecutive frames to get the answer.For example:’What is the direction of the escalator in front of me?’” 8.’resource acquisition and others’ questions should be based on the real life scenarios of blind people, and the questions must be practical (that is, questions that blind people would really ask in the current situation). Please generate the response with keys ’timestamp’, ’question’,’answer’and’questoin_type’.” The content of the key ’timestamp’ the time value corresponding to the current picture,for example:the first image corresponds to the first element in timestamp list, and so on.”The ’timestamp’ value can be found here:”f”Timestamp list: time.The first image corresponds to the first element in this list, and so on.The content of the key ’question’ is the question you generated.The content of the key ’answer’ is the answer to the corresponding question you generated.The key ’question_type’ represents which category the question type you generated belongs to.You should all generate ’resource acquisition and others’,your response should look like this example:’timestamp’:’3000’,’question’:’Is there a toilet?’,’answer’:’yes, just go straight ahead.’,’question_type’:’resource acquisition and others’.Please generate complete data, complete keys, and key values also need to be complete.Please ask questions from first person perspective.Please ask a question for each picture, and the questions asked for different pictures input at the same time need to be different.
Evaluation Prompts
You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task: ##INSTRUCTIONS: - Focus on meaningful matches: Assess whether the predicted answer and the correct answer have a meaningful match, not just literal word-for-word matches. - Criteria for Correctness: The predicted answer is considered correct if it reasonably matches any of the four standard answers, recognizing that synonyms or varied expressions that convey the same meaning are acceptable. - Allow for Paraphrasing: Understand that different wording that conveys the same fundamental idea is valid. Evaluate if the essence of the predicted answer captures the core information of the correct answer. - Flexibility in Evaluation: Use judgment to decide if variations in the predicted answer still correctly address the question, even if they do not directly replicate the correct answer’s phrasing. For example, when the correct answer is ’Left front’, Predicted Answer: ’About ten meters to your left front’, these two answers match. Please evaluate the following video-based question-answer pair: Question: {question}
Correct Answer0: {answer0}
Correct Answer1: {answer1}
Correct Answer2: {answer2}
Correct Answer3: {answer3}
Predicted Answer: {pred} Provide your evaluation only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match. Please generate the response in the form of a Python dictionary string with keys ’pred’ and ’score’, where the value of ’pred’ is a string of ’yes’ or ’no’ and the value of ’score’ is in INTEGER, not STRING. DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: {’pred’: ’yes’, ’score’: 4.8}.
Table 9: Prompts for GPT-4o mini to evaluate MLLMs on EgoBlind.